You invest NZD 8,000 in a US Vanguard ETF at a price of US$134.28. The current NZD/USD exchange rate is 0.731.

In a year, the ETF trades at $141.10, and the NZD/USD exchange rate is 0.759.

Compute your return on investment. Assume assets are divisible (i.e., you can buy a part of an ETF)
Answer #1

Amount converted = NZD amount invested*exchange rate = 8000*0.731=5848

Share bought = Amount converted/USD price = 5848/134.28=43.5507

Final amount in NZD = Share bought*final price/(NZD/USD) = 43.5507*141.1/0.759=8096.1841

Profit = Final amount in NZD-amount invested = 8096.1841-8000=96.1841

Return = profit/amount invested = 96.1841/8000=1.202%